OpenStudy (anonymous):

Please help with AP Stats!!! I will fan and medal!!! A. What’s the equation of the regression line you’d use to predict stopping distance based on speed? B. Interpret the slope of the regression line. C. Find a 95% confidence interval for the slope of the regression line. D. Test the hypothesis H0: β1 = 0 and Ha: β1 ≠ 0. E. Suppose, instead of testing H0: β1 = 0, you decided to test H0: β1 = 5 against Ha: β1 ≠ 5. What P-value would be associated with this test?

OpenStudy (anonymous):

The following image goes with the question.

OpenStudy (anonymous):

help me an I'll help you? Deal

OpenStudy (anonymous):

yeah i already learned it

OpenStudy (amistre64):

the equation is simple enough, it just takes a lot of time by hand. using point slope formula: y-Yavg = m(x-Xavg) finding the average x and y is simple enough .... the slope can be harrowing tho but theres a simple formula to gather the steps: \[m=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}\]

OpenStudy (amistre64):

we need to calculate the sums ... sum of x^2, sum of xy, sum of x, sum of y and use them in the formula

OpenStudy (anonymous):

So it that just for A?

OpenStudy (amistre64):

if the data could be copied and pasted, excel makes short work of it

OpenStudy (anonymous):

Hmm... It could be attached as a file

OpenStudy (amistre64):

or if you can type them in here directly, i can copy paste them into an excel sheet

OpenStudy (anonymous):

You just want me to copy and paste the picture I had attached?

OpenStudy (amistre64):

i cant say that im all that familiar with the CDE parts ... the data for speed and stop distance. i cant copy paste from the pictures, and its not worth the effort for me to transcribe it by hand :)

OpenStudy (anonymous):

Data Display Row Speed StopDist 1 25 63 2 25 56 3 30 84 4 35 107 5 45 153 6 45 164 7 55 204 8 55 220 9 65 285 10 65 303 Descriptive Statistics Variable N Mean Median TrMean StDev SEMean Speed 10 44.50 45.00 44.38 15.36 4.86 StopDist 10 163.9 158.5 160.0 88.4 28.0 Variable Min Max Q1 Q3 Speed 25.00 65.00 28.75 57.50 StopDist 56.0 303.0 78.7 236.2 Regression Analysis The regression equation is StopDist = __________________________________ Predictor Coef Stdev t-ratio p Constant .89.99 12.68 −7.10 0.000 Speed 5.7053 0.2707 ___________ s = 12.47 R-sq = 98.2% R-sq(adj) = 98.0% Does this help at all?

OpenStudy (amistre64):

http://stattrek.com/regression/slope-confidence-interval.aspx im reviewing this to see how to do parts C ... part B should be straight from your course material.

OpenStudy (anonymous):

This is probably easier to see and copy and paste.

OpenStudy (amistre64):

speed^2 speed stop sp*st 625 25 63 1575 625 25 56 1400 900 30 84 2520 1225 35 107 3745 2025 45 153 6885 2025 45 164 7380 3025 55 204 11220 3025 55 220 12100 4225 65 285 18525 4225 65 303 19695 21925 445 1639 85045 10 85045 445 1639 121095 5.705300353 10 21925 445 445 21225 slope

OpenStudy (anonymous):

Thank you! What letter does that refer to?

OpenStudy (amistre64):

the columns are simply the parts needed, the at the bottom of the columns are the sums of the column.... the rest is getting it into the formula and dividing it out to get a slope of 121095/21225 or simply put as 5.71 for an approximation ... how accurate it gets is up to your material a spose

OpenStudy (amistre64):

Xavg is given in the data as 44.50 Yavg is given in the data as 163.9 this creates the formula as: y = 5.71(x-44.5) - 163.9 again youll have to determine the decimal accuracy for the slope based on your materials suggestions

OpenStudy (anonymous):

Thanks for the help, so that final answer refers to part A?

OpenStudy (amistre64):

yep, and excel has a regression line function that does it all for us if need be, but i always like to do the work to keep in shape :)

OpenStudy (amistre64):

i didnt calculate the average xy parts, i assumed the data was correct as given ... might need to dbl chk

OpenStudy (amistre64):

yeah, they are good

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

gotta typo in the equation ... y = 5.71(x-44.5) + 163.9 , thats better, i had -163

OpenStudy (amistre64):

yep the stat link gives us a standard error formula to work out the confidence interval with ... SE = sb1 = sqrt [ Σ(yi - ŷi)2 / (n - 2) ] / sqrt [ Σ(xi - x)2 ]

OpenStudy (amistre64):

10180.81 380.25 11642.41 380.25 6384.01 210.25 3237.61 90.25 118.81 0.25 0.01 0.25 1608.01 110.25 3147.21 110.25 14665.21 420.25 19348.81 420.25 8791.6125 2122.5 2.035215584 is what im getting for the standard error, we use this as a standard deviation

OpenStudy (amistre64):

got that a little off, the y parts done use the average y, they use the regression line

OpenStudy (amistre64):

\[\hat y=mx+b\]is what is used to determine\[\sum(y-\hat y)^2\] not \(\bar y\)

OpenStudy (amistre64):

that seems to be a better value of the standard error, the deviation of the actual y values from the predicted y values that the line equation gives us

OpenStudy (amistre64):

the stat link says that we want the t score related the the test statistic (our equations slope), based on n-2 degrees of freedom

OpenStudy (amistre64):

any of this ringing any bells? im having to teach myself as we go for CDE

OpenStudy (amistre64):

the confidence interval is similar to previous CI calculations:\[m\pm t_{n-1}*(se)/\sqrt{(\sum(x-\bar x)^2}\]

OpenStudy (amistre64):

t_(n-2) that is

OpenStudy (anonymous):

It is making more sense. Is this referring to part C?

OpenStudy (amistre64):

hopefully :)

OpenStudy (amistre64):

yeah, finding a confidence interval

OpenStudy (amistre64):

i see the 0.2707 in the data

OpenStudy (amistre64):

i need to revise my interval notation: one website gave us the whole SE, and another im viewing split it into parts :) \[m\pm t_{n-2}*SE\]

OpenStudy (anonymous):

Okay! I don't mean to ask for too much, but before you leave, is there any way you could consolidate the final answers into A) ___, B) ____, C)____ ... etc, because I am having a hard time keeping track. Thank you so much, you have already gone above and beyond.

OpenStudy (amistre64):

B will have to come from your course material, its just the definition of what the slope tells us about the information -- something to do with correlation. A, im just now gathering from the data ... 'm' value is sitting right there near the bottom next to speed. sneaky critters, so we just use that and the averages to determine the equation with

OpenStudy (amistre64):

A, y = 5.7053x - 5.7053(44.5) + 163.9, simplify as needed C, the SE is right next to the place the slope is at .... so we need to determine the t score for a 99% CI with 10-2=8 degrees of freedom

OpenStudy (anonymous):

Oh right, I will look for B in the textbook.

OpenStudy (amistre64):

http://onlinestatbook.com/2/calculators/inverse_t_dist.html this looks useful, even tho they cant spell confidence .... 8 degrees at 99% gives us a t score of 3.355 so out confidence interval is simply\[5.7053\pm3.355(.2707)\]

OpenStudy (anonymous):

But it is asking for 95% isn't it?

OpenStudy (amistre64):

these tired old eyes of mine .... yeah, so just replace it with the appropriate t value 10-2 = 8 degrees of freedom at 95% :)

OpenStudy (amistre64):

hypothesis testing is simple enough, youll just have to see if there are any special deals when playing with the slope of the line i got no idea what we would use off hand for determing the parameters of the distribution :/ good luck with the rest of it tho.

OpenStudy (amistre64):

it looks like the test statistic for the hypothesis of Ho=0, is simply the slope divided by the SE

OpenStudy (amistre64):

\[test stat=\frac{5.7053-H_o}{.2707}\]

OpenStudy (amistre64):

the test statistic has a student t distribution, not a normal distribution .... so the P-value is the area of the 2tails associated with it.

OpenStudy (amistre64):

..and thats all i got for this :) again, good luck

OpenStudy (anonymous):

Thank you very much! I appreciate all the help.

OpenStudy (amistre64):

youre welcome